A cyclist of mass 2kg is riding his 13kg bicycle up a 1 in 10 incline. What power does he need to develop in order to maintain a constant speed of 24kph

1
Expert's answer
2022-02-18T08:22:16-0500

Given:

"m=\rm 2\: kg+13\: kg=15\: kg"

"\sin\theta=1/10"

"v=24\:\rm km/hr=6.67\: m/s"


The net force




"F=mg\sin\theta=15*9.8*1/10=14.7\:\rm N"

The total work done




"W=Fd=Fvt"


The required power



"P=\frac{W}{t}=Fv=14.7*6.67=98.0\:\rm W"